OpenStudy (anonymous):

Will Medal and fan!!! Evaluate the expression when a = –1.4 and b = –2.7. |a| – b A. 4.1 B. –1.3 C. –4.1 D. 1.3

jimthompson5910 (jim_thompson5910):

if a = -1.4, then what is |a| equal to?

OpenStudy (anonymous):

1.4?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

|a| – b = |-1.4| - (-2.7) |a| – b = 1.4 - (-2.7) |a| – b = 1.4 + 2.7 ... two negatives make a positive I'm sure you see how to finish up

OpenStudy (anonymous):

Is it 4.1 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

yw